### v1

# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

Abstract
We first remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of the same result. The first proof is a specialization of a proof of the Lindemann– Weierstrass theorem given by Beukers, Bézivin and Roba [2]. The second proof uses improper integrals of formal power series and is due to this author. At the beginning, we explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, it avoids uncountable sets. It is due to this author in [4].
1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
Hilbert’s proof [3] of the transcendence of the number $\mathrm{e}$ goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
$P:=a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0\,.$
We may assume that $a_{0}\neq 0$. For $r\in\mathbb{N}$ we define integral polynomials
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
with $n$ as in $P$. We have the following expansions starting from the lowest powers of $x$:
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+\bullet\cdot x^{r+1}+\dots$
and
$p_{r}(x+j)=\bullet\cdot x^{r+1}+\bullet\cdot x^{r+2}+\dots$
for $j=1,2,\dots,n$ and some $\bullet\in\mathbb{Z}$ (the integers). We consider integrals
$I_{r}=\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
Using additivity of the improper integral $\int_{0}^{+\infty}$ we express the product $P\cdot I_{r}$ as
$0$ $=$ $P\cdot I_{r}=A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it for $r\in\mathbb{N}$ by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r}$
where $c>1$ is a constant depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$.
The heart of the proof is the determination of the second quantity
$B_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c>1$, the above equalities
$A_{r}+B_{r}=0,\ r\in\mathbb{N}\,$,
are contradictory and we are done. In order to prove the claim, we use the integral identity that
$\int_{0}^{+\infty}x^{n}\mathrm{e}^{-x}\,\mathrm{d}x=n!$}
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). It is due to Euler and is easy to establish by induction on $n$ via integration by parts. By the linearity of the improper integral $\int_{0}^{+\infty}$, for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ we have more generally that
$\int_{0}^{+\infty}\big(\sum_{j=0}^{m}b_{j}x^{j}\big)\mathrm{e}^{-x}\,\mathrm{d}x=\sum_{j=0}^{m}b_{j}\cdot j!\,.$}
Then
$B_{r}$ $=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{j}^{+\infty}p_{r}((x-j)+j)\cdot\mathrm{e}^{-(x-j)}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{0}^{+\infty}p_{r}(y+j)\cdot\mathrm{e}^{-y}\,\mathrm{d}y}$
$=$ $\pm a_{0}(n!)^{r+1}\cdot r!+m_{r}\cdot(r+1)!\,$,
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$, in lowest powers of $x$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Interestingly, Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof that uses only proper integrals $\int_{a}^{b}$.
Hilbert’s proof is elegant and short, but it is not completely satisfactory because it employs uncountable sets; such as, for example, the integrands in Euler’s identity. We explain how to fix it and how to prove the transcendence of the number $\mathrm{e}$ so that two conditions are satisfied.
Only hereditarily at most countable (HMC) sets are directly used in the proof.
No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
$x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
the set $x_{n}$ is finite or countable. As for uses of mathematical logic, we quote a referee of [4]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$.” The reader can check [5] for $\mathrm{ACA}_{0}$. Our more demanding goal is to obtain a proof of the transcendence of $\mathrm{e}$ satisfying besides (A) also (B), that is, such that no HMC set and no mathematical logic is used.
There are (at least) three ways to achieve this goal. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
$f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
where $\mathbb{Q}$ is the (HMC) set of fractions. If real numbers are represented by Dedekind cuts, any such function $f$ is an HMC set. We realized the first way in [4]. Is there any difference in treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic.” There is an obvious meta-argument refuting the referee’s first claim: if it were the case that “everything works out in the usual way,” then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [4] for details. We include the referee’s second claim just to reiterate that our approach in [4] satisfies condition (B). In [4], we do not use technical notions of mathematical logic at all; therefore, our approach is very different from any countable treatment of real analysis offered by mathematical logic, such as, to give besides Das Kontinuum more modern references, [7] or Weaver’s forthcoming book [8].
We included this analytical digression to place the algebraic results of the present article in context. We present two algebraic proofs of the transcendence of $\mathrm{e}$ satisfying conditions (A) and (B). They are based on algebraic manipulations with formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it in early versions of [4], but we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only by mention, as “harbors” of the main players. We may say, for example, that for every two real sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots$; it is clear that $\mathbb{R}$ is used only indirectly and can be easily formulated away. We allow such uses of non-HMC sets for the sake of natural and easily understandable formulations.
In conclusion of this introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the (usually non-HMC) set of maps $f\colon\mathbb{N}_{0}\to R$. They are traditionally written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}\,,
where $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also by $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
${\textstyle\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}a_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is (formal) derivative,
$\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
So for derivatives the characteristic of $R$ becomes relevant. Less often one uses the (formal) primitive
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,$,}
but in this article, it is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$ that satisfy
$\lim_{n\to\infty}a_{n}c^{n}=0$
for every $c\in K$. In other words, $f(x)$ has an infinite radius of convergence.
The binary and unary operations on $R[[x]]$ that we have listed so far, ++, $\cdot$, $(\dots)^{\prime}$, and $\int(\dots)$, are formal in the sense that, for example, $x_{n}](f\cdot g)$ is determined by finitely many coefficients in $f$ and $g$, and similarly for the other operations. In formal arguments involving elements $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$ it is also forbidden to use values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k}$}
of $f(x)$ on elements $c\in K$. It is still an open problem to find a fully formal proof of the transcendence of $\mathrm{e}$ based on formal power series; or at least no such proof is known to this author. However, in the next two sections we present two such proofs if formality is relaxed to quasiformality.
An argument involving elements of $K[[x]]$ is quasiformal if it uses only formal and quasiformal operations, and for every $f(x)\in K[[x]]$ appearing in it only at most countably many values $f(c)$ for $c\in K$ are used. Quasiformal operations allow uses of limit transitions in $K$ and of at most countably many values $f(c)$.
Quasiformal operations on $\mathbb{R}[[x]]$ are introduced, for example, in Definitions 3.3 and 3.11.
2 Specializing the proof of Beukers, Bézivin and Robba
The Lindemann–Weierstrass theorem says that
$b_{1}\mathrm{e}^{a_{1}}+b_{2}\mathrm{e}^{a_{2}}+\dots+b_{t}\mathrm{e}^{a_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin and Robba gave in [2] another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly describe logical structure of the proof, we completely reorganize the exposition in [2]. At the end of this section, we explain why this proof is semiformal (and not formal), and why it satisfies conditions (A) and (B).
We begin by recalling that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition in partial fractions
$f(x)=r(x)+\sum_{j=1}^{t}\sum_{i=1}^{m_{j}}\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}\,$,}
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$ and the $\alpha_{j}$ are nonzero and mutually distinct; for $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
$q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
The partial fractions are the formal power series
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\beta_{j,i}\sum_{n\geq 0}\binom{-i}{n}\alpha_{j}^{n}x^{n}\,.$}
We say that $1/\alpha_{j}$ is a pole of $f(x)$ with order $m_{j}$.
Proposition 2.1
If $p(x)\in\mathbb{C}[x]$, $m\in\mathbb{N}_{0}$, $c\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$) and if $A(x)$ in $\mathbb{C}[[x]]$ is rational, then the rational formal power series
$p(x)A(x)+cx^{m}A^{\prime}(x)$
either has no pole (is a polynomial) or all its poles have orders at least $2$.
Proof. This follows from the definition of the decomposition of a rational formal power series in partial fractions. The product $p(x)A(x)$ possibly only decreases the orders of the poles of $A(x)$, or cancels them and does not create any new poles. The product $cx^{m}A^{\prime}(x)$ preserves all poles of $A(x)$, increases their orders by $1$, and does not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. Therefore, the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We prove that $\mathrm{e}$ is transcendental by means of the following theorem on rationality of formal power series, which is extracted from the arguments in [2].
Theorem 2.2
Let $t\in\mathbb{N}$, $b_{1}$, $b_{2}$, $\dots$, $b_{t}$ be nonzero integers and $\alpha_{1}$, $\alpha_{2}$, $\dots$, $\alpha_{t}$ be distinct natural numbers. For every $n\in\mathbb{N}_{0}$ we define integers
$u_{n}=b_{1}\alpha_{1}^{n}+\dots+b_{t}\alpha_{t}^{n}\,\text{ and }\,v_{n}=n!\cdot{\textstyle\sum_{r=0}^{n}u_{r}/r!}\,.$
Suppose that for some constant $A\geq 1$ we have bound
$v_{n}=O(A^{n})\,\text{ for }\,n\in\mathbb{N}_{0}\,.$
Then the formal power series $\sum_{n\geq 0}v_{n}x^{n}$ is rational, because
$\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\in\mathbb{Z}[x]$}
for some $k\in\mathbb{N}$, where the integers $a_{j}$ are the coefficients in
$\prod_{j=1}^{t}(1-\alpha_{j}x)=1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\,.$}
Proof. For $k,n\in\mathbb{N}_{0}$ we define integers $v_{n}(k)$ by the relations
$\sum_{n\geq 0}v_{n}(k)x^{n}=\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\,.$}
Then for every $n\geq t$ and $k\in\mathbb{N}_{0}$ we have the recurrence
$v_{n}(k+1)=v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)\,.$
We establish for $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
  1. 1.
    We have bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c>0$ is a constant.
  2. 2.
    $k!\,|\,v_{n}(k)$.
The bound in item 1 follows by induction on $k$ and the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound clearly holds. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
$|v_{n}(k+1)|$ $=$ $|v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)|$
$\leq$ $c\big(A^{n}C^{k}+A^{n-1}C^{k}|a_{1}|+\dots+A^{n-t}C^{k}|a_{t}|\big)\leq cA^{n}C^{k+1}\,.$
In order to prove the divisibility in item 2, we note that $v_{0}=u_{0}$ and that for $n\in\mathbb{N}$,
$\frac{v_{n}}{n!}-\frac{v_{n-1}}{(n-1)!}=\frac{u_{n}}{n!}\,\text{ (hence $v_{n}-nv_{n-1}=u_{n}$)}\,.$}
For $k=0$ the divisibility trivially holds. So let $k\in\mathbb{N}$. We set $(0)_{0}=(n)_{0}=1$ and $(n)_{m}=n(n-1)\dots(n-m+1)$ for $m\in\mathbb{N}$. For $n\in\mathbb{N}_{0}$ we combine $\min(n,k)$ displayed differences, multiply the result by $n!$, and get that
$v_{n}=(n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}+(n)_{k}v_{n-k}$}\,,
where the $u_{\dots}$ and $v_{\dots}$ with negative indices are defined arbitrarily, for example as 0. We denote the last summand $(n)_{k}v_{n-k}$ ($\in\mathbb{Z}$) by $w_{n}$. Note that always $k!\,|\,w_{n}$ because $w_{n}=k!\cdot\binom{n}{k}v_{n-k}$ for $n\geq k$, and $w_{n}=0$ for $0\leq n<k$. The difference of the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ is
$V(x)-W(x)=\sum_{n\geq 0}\big((n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}\big)x^{n}\,.$}
We claim that
$V(x)-W(x)=\frac{p(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{k}}=:\frac{p(x)}{q(x)}$}
for some polynomial $p(x)\in\mathbb{Z}[x]$ with degree less than $tk$. Then for every $n\geq tk$ we have
$v_{n}(k)=[x^{n}]q(x)V(x)=[x^{n}]q(x)W(x)$
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for any $r$ with $0\leq r\leq k-1$ we indeed have, as $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
${\textstyle\sum_{n\geq 0}(n)_{r}u_{n-r}x^{n}}$ $=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}\sum_{n\geq r}\binom{n}{r}\alpha_{j}^{n-r}x^{n-r}}$
$=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}(1-\alpha_{j}x)^{-r-1}=\frac{p_{r}(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{r+1}}}$
for some polynomial $p_{r}(x)\in\mathbb{Z}[x]$ with degree less than $t(r+1)$.
Now the two properties of numbers $v_{n}(k)$ imply that if $n\geq tk$ and $v_{n}(k)\neq 0$, then
$k!\leq|v_{n}(k)|\leq cA^{n}C^{k}\,.$
In other words, if $n\geq tk$ and if $k!>cA^{n}C^{k}$, then $v_{n}(k)=0$. Using this implication and the modification
$v_{n}(k)=v_{n}(k+1)+a_{1}v_{n-1}(k)+\dots+a_{t}v_{n-t}(k),\ n\geq t\,$,
of the above recurrence, we show that for a fixed $k\in\mathbb{N}$ we have $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
$N$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge tk\leq n\leq 2tk\}\,\text{ and}$
$M$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge n\geq 2tk\}\,.$
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the ordered pairs in the modified recurrence:
$\gamma=\langle k,\,n\rangle,\,\beta_{0}=\langle k+1,\,n\rangle,\,\beta_{1}=\langle k,\,n-1\rangle,\,\dots,\,\beta_{t}=\langle k,\,n-t\rangle\,.$
The following three properties of the norm $\|\dots\|$ and the pairs $\gamma$, $\beta_{0}$, $\dots$, $\beta_{t}$ are easy to prove.
  1. 1.
    The norm $\|\dots\|$ attains a minimum value on the set $M$ (namely, zero).
  2. 2.
    If $\gamma\in M$ then $\beta_{i}\in M\cup N$ for every $i=0,1,\dots,t$.
  3. 3.
    For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
From these three properties and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every $k\geq k_{0}$ has the required property. $\Box$
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j\in[t]$ ($\{1,2,\dots,t\}$), are $2t$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
$\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
We may assume that $\alpha_{j}\in\mathbb{N}$. We set $A:=\max(\{\alpha_{1},\dots,\alpha_{t}\})$ and define $u_{n}$ and $v_{n}$ as in Theorem 2.2. Then
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
and
$\sum_{n\geq 0}\frac{u_{n}}{n!}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
From the definition of $u_{n}$ and $A$ it easily follows that $u_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Since by the definition of $v_{n}$ we have
$|v_{n}|={\textstyle n!\cdot\big|\sum_{r=0}^{n}u_{r}/r!\big|=n!\cdot\big|\sum_{r=n+1}^{\infty}u_{r}/r!\big|}$
— now we invoke the contradictory assumption — we get the bound
$|v_{n}|$ $\leq$ ${\textstyle\frac{|u_{n+1}|}{n+1}+\frac{|u_{n+2}|}{(n+1)(n+2)}+\dots\leq cA^{n}\big(\frac{A}{n+1}+\frac{A^{2}}{(n+1)(n+2)}+\dots\big)}$
$\leq$ ${\textstyle cA^{n}\sum_{j\geq 0}\frac{A^{j}}{j!}=c\,\mathrm{e}^{A}\cdot A^{n}\,,}$
where $c>0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{n\geq 0}(v_{n}-nv_{n-1})x^{n}=\sum_{n\geq 0}u_{n}x^{n}=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
(with $v_{-1}:=0$). However, the equality
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
We reiterate that the proof in [2] is more general as it establishes the full LW theorem. In this situation, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is only semiformal and not formal because the key player is the infinite sum
$\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal feature of it.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not directly used in the proof. Condition (B), which states that no formalism of mathematical logic is used, is clearly satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem) that satisfies conditions (A) and (B), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
3 The proof by formal integration
We present the second semiformal proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [4]. The idea is simply to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series. At the end of the section we explain why the proof is semiformal and why conditions (A) and (B) are satisfied.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce additional operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we reviewed in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
$f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
This is a formal operation. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
$(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$.
In Section 1 we defined the derivative $f^{\prime}$ and the primitive $\int f$ of a formal power series $f\in\mathbb{R}[[x]]$. In the next proposition, whose straightforward proof we omit, we collect properties of these operations.
Proposition 3.1
Let $f,g\in\mathbb{R}[[x]]$ and $a,b\in\mathbb{R}$. The following holds.
  1. 1.
    $(af+bg)^{\prime}=af^{\prime}+bg^{\prime}$, $(fg)^{\prime}=f^{\prime}g+fg^{\prime}$ and $f(ax)^{\prime}=af^{\prime}(ax)$.
  2. 2.
    $\int(af+bg)=a\int f+b\int g$ and $\int f(ax)=\frac{1}{a}(\int f)(ax)$ if $a\neq 0$.
  3. 3.
    If $f\in\mathbb{R}\{x\}$ then $f^{\prime},\int f\in\mathbb{R}\{x\}$.
  4. 4.
    $\big(\int f\big)^{\prime}=f$ and $\int(f^{\prime})=f-[x^{0}]f$.
We define the (formal) exponential.
Definition 3.2
The formal power series
$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
is called the (formal) exponential.
It is easy to show that $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and that $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a semiformal operation.
Definition 3.3 (semiformal shifts)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ be in $\mathbb{R}\{x\}$ and let $b\in\mathbb{R}$. We define the formal power series $f(x+b)$ in $\mathbb{R}[[x]]$ by
$f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent, and $f(x+b)\in\mathbb{R}\{x\}$. It is easy to see that $f(x+0)=f(x)$ and that if $f(x)\in\mathbb{R}[x]$ then $f(x+b)$ is the usual substitution in a polynomial. In the next two propositions we show that semiformal shifts behave in the expected way.
Proposition 3.4
Let $f(x)\in\mathbb{R}\{x\}$, $c,d\in\mathbb{R}$ and $g(x)=f(x+c)$. Then
g(x+d)=f(x+(c+d)).g(x+d)=f(x+(c+d))\;.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $c$, $d$ and $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ be as stated. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,g(x+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus the coefficient of $x^{n}$ in $g(x+d)$ equals
${\textstyle\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
$={\textstyle\sum_{m\geq n}\sum_{l\geq m}\frac{l!}{n!(m-n)!(l-m)!}a_{l}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}\,,}$
which is the coefficient of $x^{n}$ in $f(x+(c+d))$. In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
Proposition 3.5
Let $f,g\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
$(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus the coefficient of $x^{n}$ in the product $f(x+c)\cdot g(x+c)$ equals
${\textstyle\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\cdot\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)=}$
${\textstyle=\sum_{k=0}^{n}\sum_{m\geq k}\sum_{l\geq n-k}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}\,,}$
which is the coefficient of $x^{n}$ in $(f\cdot g)(x+c)$. In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([6]). $\Box$
We prove a semiformal exponential identity.
Theorem 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$ and $f(x)=\mathrm{Exp}\big(x+\frac{b}{a}\big)$. Then
$f(ax)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
Proof. Let $a$, $b$ and $f(x)$ be as stated. Then, indeed,
$f(ax)$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)\frac{a^{n}}{n!}x^{n}}$
$=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\;.$
$\Box$
We introduce the Newton integral of convergent formal power series.
Definition 3.7 (Newton integral of fps)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
${\textstyle\int_{u}^{v}f=\int_{u}^{v}f(x)}$ $:=$ ${\textstyle\big(\int f\big)(v)-\big(\int f\big)(u)}$
$=$ ${\textstyle\sum_{n\geq 0}a_{n}\cdot\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}a_{n}\cdot\frac{u^{n+1}}{n+1}}$
of the values of the formal primitive of $f$ at the numbers $v$ and $u$.
The integral is always defined by item 3 of Proposition 3.1 because $\int f\in\mathbb{R}\{x\}$. The Newton integral is a semiformal operation because it uses two values of a formal power series. We establish for it three integral transformations: linear combination, additive combination, and shift of the integration interval. The first two are quite straightforward.
Proposition 3.8
Let $f,g\in\mathbb{R}\{x\}$ and $u,v,a,b\in\mathbb{R}$. Then
$\int_{u}^{v}(af+bg)=a\int_{u}^{v}f+b\int_{u}^{v}g\,.$}
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in Proposition 3.1. $\Box$
Proposition 3.9
Let $f\in\mathbb{R}\{x\}$ and $u,v,w\in\mathbb{R}$. Then
$\int_{u}^{w}f=\int_{u}^{v}f+\int_{v}^{w}f\,.$}
Proof. Indeed,
$\big(\int f\big)(w)-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+\big(\int f\big)(w)-\big(\int f\big)(v)\,.$}
$\Box$
The third transformation needs some work.
Proposition 3.10
Let $f\in\mathbb{R}\{x\}$ and $u,v,b\in\mathbb{R}$. Then
$\int_{u}^{v}f(x+b)=\int_{u+b}^{v+b}f(x)\,.$}
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$, $v$ and $b$ be as stated. Then
${\textstyle\int_{u}^{v}f(x+b)=}$
${\textstyle=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{u^{n+1}}{n+1}}$
$={\textstyle\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}v^{n+1}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}u^{n+1}}$
${\textstyle=\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}v^{l}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}u^{l}}$
${\textstyle=\sum_{m\geq 0}a_{m}\frac{(v+b)^{m+1}}{m+1}-\sum_{m\geq 0}a_{m}\frac{(u+b)^{m+1}}{m+1}=\int_{u+b}^{v+b}f(x)\,.}$
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add and subtract the term $b^{m+1}$ in the two inner sums. In the fifth line, we use the binomial theorem. $\Box$
We introduce the improper Newton integral of formal power series.
Definition 3.11 (improper Newton $\int$ of fps)
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the limit
$L=\lim_{n\to\infty}\int_{u}^{b_{n}}f\ \ (\in\mathbb{R})$}
exists, then $L$ does not depend on the sequence $(b_{n})$ and we define
$\int_{u}^{+\infty}f=\int_{u}^{+\infty}f(x):=L\,.$}
We call this integral the improper Newton integral from $u$ to $\infty$ of the formal power series $f(x)$.
We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
We obtain for the improper integral the same three integral transformations as before.
Proposition 3.12
Let $f,g\in\mathbb{R}\{x\}$ and let $u,a,b\in\mathbb{R}$. Then the equality
$\int_{u}^{+\infty}(af+bg)=a\int_{u}^{+\infty}f+b\int_{u}^{+\infty}g$}
holds whenever the last two improper Newton integrals exist.
Proof. This follows from Proposition 3.8 and the arithmetic of limits of real sequences. $\Box$
Proposition 3.13
Let $f\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f=\int_{u}^{v}f+\int_{v}^{+\infty}f$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ exists. But then
${\textstyle\int_{u}^{+\infty}f=}$
${\textstyle=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
The third transformation is again less immediate.
Proposition 3.14
Let $f\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f(x+b)=\int_{u+b}^{+\infty}f(x)$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$ and $b$ be as stated. Let $(c_{n})$ be any real sequence converging to $\infty$, and let $L$ and $R$ be the integrals on the left-hand side and the right-hand side, respectively. It suffices to show the implication $\exists\,R\Rightarrow R=L$; the opposite implication then follows. Indeed, we suppose that integral $L$ exists and write it as
$L=\int_{u+b-b}^{+\infty}f(x+b)\,.$}
We set $g(x)=f(x+b)$, use the assumed implication with b-b and get that
[…]
Thus, the implication $\exists\,L\Rightarrow L=R$ holds as well.
We suppose that integral $R$ exists. We replace the test sequence $(c_{n})$ with $(c_{n}+b)$ and get
$R$ $=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{n\geq 0}a_{n}\frac{(c_{m}+b)^{n+1}}{n+1}\big)-\sum_{n\geq 0}a_{n}\frac{(u+b)^{n+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{n+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n+1}{l+1}a_{n}b^{n-l}\frac{u^{l+1}}{n+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{c_{m}^{l+1}}{l+1}\big)-\sum_{l\geq 0}\sum_{n\geq l}\binom{n}{l}a_{n}b^{n-l}\frac{u^{l+1}}{l+1}}$
$=$ $\lim_{m\to\infty}{\textstyle\big(\int f(x+b)\big)(c_{m})}-{\textstyle\big(\int f(x+b)\big)(u)}=L\,.$
In the second line, we used the binomial theorem, and we changed the order of summation, which is allowed by the convergence of $f(x)$. Also, we canceled the summands corresponding to $l+1=0$. In the third line, we used the transformation $\binom{l+1}{n+1}=\frac{l+1}{n+1}\binom{l}{n}$. In the fourth line, we use the definition of formal primitive and Definition 3.3. $\Box$
We need bounds on integrals
$\int_{0}^{i}p(x)\mathrm{Exp}(-x)$}
where $i\in\mathbb{N}_{0}$, $p(x)$ is an integral polynomial and $\mathrm{Exp}(-x)=\mathrm{Exp}((-1)x)$. In the semiformal approach, this appears problematic because one deduces the standard LM bound by considering all values of the integrand on $[0,i]$ or $0,i]\cap\mathbb{Q}$. Instead, we use the next substitute LM bound, which can be proven by means of just two values of the primitive.
Proposition 3.15
For every $i,k\in\mathbb{N}_{0}$,
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
Proof. The integrand
$\text{$\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus the Newton integral is in absolute value at most $\sum_{n=0}^{\infty}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}$. $\Box$
It is important that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ was Euler’s identity. Now we prove it in the setting of the improper Newton integral of formal power series.
Theorem 3.16
For every $k\in\mathbb{N}_{0}$, the integral
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n=k}^{\infty}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. For $k=0$ we use that
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,.$}
Hence
$I_{0}=\lim_{n\to\infty}{\textstyle(1-\mathrm{Exp}(-a_{n}))-(1-\mathrm{Exp}(-0))=1-0=1\,.}$
For $k\in\mathbb{N}_{0}$, let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$. Let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}\,.$
By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}$. Thus by Proposition 3.12 and induction,
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}=k\cdot}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}F_{k}(a_{n})+F_{k}(0)$
$=$ $k\cdot(k-1)!-0+0=k!\;.$
$\Box$
We are ready to translate Hilbert’s proof of the transcendence of $\mathrm{e}$ into formal power series. We again assume for the contradiction that
$a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral (of formal power series)
$\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)\,$,}
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ (as in Section 1). This integral exists by Proposition 3.12 and Theorem 3.16. Using Proposition 3.13 we get that
$0$ $=$ $A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)+\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
for every $r\in\mathbb{N}$. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded.
The determination of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that we calculate semiformally. We define abbreviations
$\mathrm{Exp}(-(x-i))$ $=$ $f((-1)x),\ f(x)=\mathrm{Exp}(x+(-i))\,\text{ and}$
$\mathrm{Exp}(-((x-i)+i))$ $=$ $f((-1)x),\ f(x)=g(x+i),\ g(x)=\mathrm{Exp}(x+(-i))\,.$
Then
$B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
[…]
Recall that $x$ is a formal variable. Using Proposition 3.12, Theorem 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
It remains to bound the numbers
$A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
The polynomial
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$
has degree less than $(n+1)(r+1)$ and each coefficient of it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
$(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get the same contradiction as in Section 1. $\Box$
The previous proof is quasiformal and not formal because of Definitions 3.3 and 3.11. It clearly satisfies conditions (A) and (B).