### v3

# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Then, using formal power series, we give two algebraic semiformal proofs of this result. The first proof is a specialization of the proof of the Lindemann–Weierstrass theorem found by Beukers, Bézivin and Robba [2]. The second proof uses improper integrals of formal power series and is due to this author. We explain what “semiformal” means and indicate yet another analytical proof of the transcendence of $\mathrm{e}$. Like the two algebraic proofs, this proof avoids uncountable sets. It is due to this author in [7].
1 Hilbert’s proof that $\mathrm{e}$ is transcendental. Why this proof is not completely satisfactory. How to fix it
The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4]. Hilbert’s proof [5] goes as follows. We suppose for the contradiction that there exists $n+1$, $n\in\mathbb{N}$ ($\{1,2,\dots\}$), integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
$P:=a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0\,.$
We may assume that $a_{0}\neq 0$. For $r\in\mathbb{N}$ we define integral polynomials
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
with $n$ as in $P$. We have the expansions
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+ax^{r+1}+\dots$
and
$p_{r}(x+j)=bx^{r+1}+cx^{r+2}+\dots\,$,
where $j=1,2,\dots,n$ and $a,b,c,\dots$ are in $\mathbb{Z}$ (the integers). We consider improper integrals
$I_{r}=\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
Using the additivity of improper integrals, we express the product $PI_{r}$ as
$0$ $=$ $PI_{r}=A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it for $r\in\mathbb{N}$ by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r}$
where $c\geq 1$ is a constant (depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$).
We determine the second quantity
$B_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c\geq 1$, the above equalities
$A_{r}+B_{r}=0,\ r\in\mathbb{N}\,$,
are contradictory and we are done. We prove the claim by means of the integral identity that
$\int_{0}^{+\infty}x^{n}\mathrm{e}^{-x}\,\mathrm{d}x=n!$}
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). The identity is due to Euler and is easy to establish by induction on $n$ via integration by parts. Using the linearity of improper integrals, we obtain for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ the more general identity that
$\int_{0}^{+\infty}\big(\sum_{j=0}^{m}b_{j}x^{j}\big)\mathrm{e}^{-x}\,\mathrm{d}x=\sum_{j=0}^{m}b_{j}\cdot j!\,.$}
Thus
$B_{r}$ $=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{j}^{+\infty}p_{r}((x-j)+j)\cdot\mathrm{e}^{-(x-j)}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{0}^{+\infty}p_{r}(y+j)\cdot\mathrm{e}^{-y}\,\mathrm{d}y}$
$=$ $\pm a_{0}(n!)^{r+1}\cdot r!+m_{r}\cdot(r+1)!\,$,
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof is given in Conrad [3].
Hilbert’s proof is elegant and short, but it has the unsatisfactory feature that it employs uncountable sets, for example, the integrands in Euler’s identity. We explain how to improve upon it by proving the transcendence of $\mathrm{e}$ so that two conditions hold.
Only hereditarily at most countable (HMC) sets are directly used in the proof.
No technical notion of mathematical logic such as, for example, the notion of a formula, is used in the proof.
We call proofs satisfying these two conditions AB proofs. A set $x$ is HMC if for every index $n\in\mathbb{N}_{0}$ and every chain of sets
$x_{n}\in x_{n-1}\in\dots\in x_{0}=x\,$,
the set $x_{n}$ is at most countable (finite or countable). As for uses of mathematical logic in analysis, we quote a referee of [7]: “That basic real analysis is essentially a countable endeavor has been well-known for over 100 years. My secondhand understanding is that Weyl articulated this very well in Das Kontinuum, which nowadays we would say developed analysis in $\mathrm{ACA}_{0}$”. The reader can check [8] for $\mathrm{ACA}_{0}$. Our more ambitious and demanding goal is to obtain an AB proof of the transcendence of $\mathrm{e}$. In a related result, Kaibkhanov and Skopenkov [6] gave an AB proof of the transcendence of the Mahler number $\sum_{n\geq 0}2^{-2^{n}}$ based on binary expansions.
There are at least three ways to such proof. The first way is to replace real functions of the type $f\colon[a,b]\to\mathbb{R}$, where $a<b$ are real numbers and $\mathbb{R}$ is the set of real numbers, with the functions
$f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}\,$,
where $\mathbb{Q}$ is the set of fractions (if real numbers are Dedekind cuts, every such function $f$ is an HMC set), and to develop a sufficient fragment of real analysis for such functions. We realized this way in [7]. Is there any difference in the treatment of the two kinds of functions? The above mentioned referee thinks not: “Continuous functions on the reals are determined by their values on the rationals, so everything works out in the usual way. I don’t see how what the article is doing is so very different from developing analysis in (full) second-order arithmetic”. An obvious meta-argument refutes the first claim: if it were the case that “everything works out in the usual way”, then some enthusiasts would be teaching courses in analysis based on functions of the type $f\colon[a,b]\cap\mathbb{Q}\to\mathbb{R}$ for years. Since you have not heard about such courses, there must be some difficulties with this approach. Indeed, there are, but they can be overcome. The interested reader is referred to [7] for details. The second claim is refuted by the fact that in [7] we provide an AB proof. We do not use any technical notion of mathematical logic, and our approach in [7] is therefore very different from the countable treatments of real analysis offered by mathematical logic. For examples of the latter approaches, see Weaver [10, 11].
This analytical digression places the algebraic results of the present article in context. These results are two AB proofs of the transcendence of $\mathrm{e}$. Both are based on formal power series. The first proof in Section 2 is a specialization of the proof in [2]. The second proof in Section 3 is due to this author. We presented it already in early versions of [7], and we rework it here.
The word “directly” in condition (A) means that non-HMC sets may appear in the proof only as supersets of the main, directly used sets, which have to be HMC. We may say, for example, that for every two sequences $(a_{n}),(b_{n})\subset\mathbb{R}$ we have $\dots\;$. It is clear that the set $\mathbb{R}$ is mentioned only for convenience and can be easily formulated away. We allow such uses of non-HMC sets in AB proofs for the sake of natural and easily legible formulations.
In conclusion of the introduction, we review formal power series and explain the meaning of the adjective “semiformal”. If $R$ is an (integral) domain, we write $R[[x]]$ for the uncountable set of maps $f\colon\mathbb{N}_{0}\to R$. They are written as infinite formal sums
$f=f(x)=\sum_{n\geq 0}a_{n}x^{n}$}
with $a_{n}=f(n)$. We denote the coefficient $a_{n}$ also as $x^{n}]f$. The set $R[[x]]$ is endowed with the operations
${\textstyle f(x)+g(x)=\sum_{n\geq 0}a_{n}x^{n}+\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}(a_{n}+b_{n})x^{n}\,\text{ and}}$
${\textstyle f(x)\cdot g(x)=\sum_{n\geq 0}a_{n}x^{n}\cdot\sum_{n\geq 0}b_{n}x^{n}}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{k=0}^{n}a_{k}b_{n-k}\big)x^{n}\,,}$
and neutral elements $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\dots\;$. Then
R[[x]]=R[[x]], $\langle R[[x]],\,0,\,1$$,\cdot\rangle$
is a domain and $f$ is an $R[[x]]$-unit iff $x^{0}]f$ is an $R$-unit. An important unary operation on $R[[x]]$ is the (formal) derivative,
$\big(\sum_{n\geq 0}a_{n}x^{n}\big)^{\prime}=\sum_{n\geq 0}(n+1)_{R}\cdot a_{n+1}x^{n}\,.$}
Less often one uses the (formal) primitive
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
In this article, $\int$ is more important than the derivative. If $K$ is the normed field $K=\mathbb{R}$ or $K=\mathbb{C}$ (complex numbers), we denote by $K\{x\}$ the subdomain of $K[[x]]$ formed by the convergent formal power series $f(x)=\sum_{n\geq 0}a_{n}x_{n}$. They satisfy for every $c\in K$ that
$\lim_{n\to\infty}a_{n}c^{n}=0_{K}\,.$
In other words, $f(x)\in K\{x\}$ iff $f(x)$ has an infinite radius of convergence.
The binary and unary operations on $R[[x]]$ listed so far, ++, $\cdot$, $(\cdot)^{\prime}$, and $\int(\cdot)$, are formal in the sense that, for example, for every $n\in\mathbb{N}_{0}$ the coefficient
$x^{n}](f\cdot g)$
is determined by only finitely many coefficients in $f$ and $g$, and the same holds for the other operations. In formal proofs involving formal power series $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $K[[x]]$, only formal operations may be used, and it is forbidden to use the functional values
$f(c)=\sum_{n\geq 0}a_{n}c^{n}=\lim_{n\to\infty}\sum_{k=0}^{n}a_{k}c^{k},\ c\in K\,.$}
It is an open problem to give a completely formal AB proof of the transcendence of $\mathrm{e}$ based on formal power series. No such proof is known to this author.
However, we can present two such proofs in the next two sections if formality is relaxed to semiformality. Semiformal operations (in Section 3 we introduce several of them) are defined on $K\{x\}$, may use limit transitions, and may use at most countably many values $f(c)$ for $f\in K[[x]]$ and $c\in K$. Semiformal proofs concerning elements in $K[[x]]$ are proofs that satisfy the following condition.
Only formal and semiformal operations on $K[[x]]$ are used. For every formal power series involved in the proof, only at most countably many of its values are used.
In Sections 2 and 3 we give two AB &\& SF proofs of the transcendence of $\mathrm{e}$.
2 Specializing the proof of Beukers, Bézivin and Robba
The Lindemann–Weierstrass theorem says that
$b_{1}\mathrm{e}^{\alpha_{1}}+b_{2}\mathrm{e}^{\alpha_{2}}+\dots+b_{t}\mathrm{e}^{\alpha_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t$, with $t\in\mathbb{N}$, algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly present the logical structure of their proof, we reorganize the exposition in [2].
We first recall that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational formal power series $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition into partial fractions
$f(x)=r(x)+\sum_{j=1}^{t}\sum_{i=1}^{m_{j}}\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}\,$,}
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct. For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
$q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
The partial fractions are the formal power series
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\sum_{n\geq 0}\binom{-i}{n}\beta_{j,i}\alpha_{j}^{n}\cdot x^{n}\,.$}
We say that $1/\alpha_{j}$ is a pole of $f(x)$ with order $m_{j}$.
Proposition 2.1
If $p(x)\in\mathbb{C}[x]$, $m\in\mathbb{N}_{0}$, $c\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$) and if $A(x)$ in $\mathbb{C}[[x]]$ is rational, then the rational formal power series
$p(x)A(x)+cx^{m}A^{\prime}(x)$
either has no pole (is a polynomial) or all its poles have orders at least $2$.
Proof. The product $p(x)A(x)$ only possibly decreases the orders of the poles of $A(x)$ or cancels them and does not create any new poles. The product and derivative $cx^{m}A^{\prime}(x)$ preserve all poles of $A(x)$, increase their orders by $1$, and do not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. It follows that the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We extracted the following theorem on rational formal power series from the arguments in [2].
Theorem 2.2
Let $t\in\mathbb{N}$, $b_{1}$, $b_{2}$, $\dots$, $b_{t}$ be nonzero integers and $\alpha_{1}$, $\alpha_{2}$, $\dots$, $\alpha_{t}$ be distinct natural numbers. For every $n\in\mathbb{N}_{0}$ we define integers
$u_{n}=b_{1}\alpha_{1}^{n}+\dots+b_{t}\alpha_{t}^{n}\,\text{ and }\,v_{n}=n!\cdot{\textstyle\sum_{r=0}^{n}u_{r}/r!}\,.$
Suppose that for some constant $A\geq 1$ we have bound
$v_{n}=O(A^{n})\,\text{ for }\,n\in\mathbb{N}_{0}\,.$
Then the formal power series $\sum_{n\geq 0}v_{n}x^{n}$ is rational, because
$\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\in\mathbb{Z}[x]$}
for some $k\in\mathbb{N}$, where the integers $a_{j}$ are the coefficients in
$\prod_{j=1}^{t}(1-\alpha_{j}x)=1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\,.$}
Proof. For $k,n\in\mathbb{N}_{0}$ we define integers $v_{n}(k)$ by the relations
$\sum_{n\geq 0}v_{n}(k)x^{n}=\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\,.$}
Then for every $n\geq t$ and $k\in\mathbb{N}_{0}$ we have the recurrence
$v_{n}(k+1)=v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)\,.$
We establish for every $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
  1. 1.
    We have the bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c\geq 0$ is a constant.
  2. 2.
    $k!\,|\,v_{n}(k)$.
The bound in item 1 follows by induction on $k$ from the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound holds because it is the assumed bound on $v_{n}$. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
$|v_{n}(k+1)|$ $=$ $|v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)|$
$\leq$ $c\big(A^{n}C^{k}+A^{n-1}C^{k}|a_{1}|+\dots+A^{n-t}C^{k}|a_{t}|\big)\leq cA^{n}C^{k+1}\,.$
In order to prove the divisibility in item 2, we note that $v_{0}=u_{0}$ and that for $n\in\mathbb{N}$,
$\frac{v_{n}}{n!}-\frac{v_{n-1}}{(n-1)!}=\frac{u_{n}}{n!}\,\text{ (hence $v_{n}-nv_{n-1}=u_{n}$)}\,.$}
For $k=0$ the divisibility trivially holds. So let $k\in\mathbb{N}$. We set $(0)_{0}=(n)_{0}=1$ and $(n)_{m}=n(n-1)\dots(n-m+1)$ for $m\in\mathbb{N}$. For $n\in\mathbb{N}_{0}$ we combine $\min(n,k)$ displayed differences, multiply the result by $n!$, and get that
$v_{n}=(n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}+(n)_{k}v_{n-k}$}\,,
where the $u_{\dots}$ and $v_{\dots}$ with negative indices are defined arbitrarily, for example as 0. We denote the last summand $(n)_{k}v_{n-k}$ ($\in\mathbb{Z}$) by $w_{n}$. Note that always $k!\,|\,w_{n}$ because $w_{n}=k!\cdot\binom{n}{k}v_{n-k}$ for $n\geq k$, and $w_{n}=0$ for $0\leq n<k$. The difference of the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ is
$V(x)-W(x)=\sum_{n\geq 0}\big((n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}\big)x^{n}\,.$}
We claim that
$V(x)-W(x)=\frac{p(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{k}}=:\frac{p(x)}{q(x)}$}
for some polynomial $p(x)\in\mathbb{Z}[x]$ with degree less than $tk$. Then for every $n\geq tk$ we have
$v_{n}(k)=[x^{n}]q(x)V(x)=[x^{n}]q(x)W(x)$
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for any $r$ with $0\leq r\leq k-1$ we indeed have, as $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
${\textstyle\sum_{n\geq 0}(n)_{r}u_{n-r}x^{n}}$ $=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}\sum_{n\geq r}\binom{n}{r}\alpha_{j}^{n-r}x^{n-r}}$
$=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}(1-\alpha_{j}x)^{-r-1}=\frac{p_{r}(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{r+1}}}$
for some polynomial $p_{r}(x)\in\mathbb{Z}[x]$ with degree less than $t(r+1)$.
Now the two properties of numbers $v_{n}(k)$ imply that if $n\geq tk$ and $v_{n}(k)\neq 0$, then
$k!\leq|v_{n}(k)|\leq cA^{n}C^{k}\,.$
In other words, if $n\geq tk$ and if $k!>cA^{n}C^{k}$, then $v_{n}(k)=0$. Using this implication and the modification
$v_{n}(k)=v_{n}(k+1)+a_{1}v_{n-1}(k)+\dots+a_{t}v_{n-t}(k),\ n\geq t\,$,
of the above recurrence, we show that there exists a $k\in\mathbb{N}$ such that $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
$N$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge tk\leq n\leq 2tk\}\,\text{ and}$
$M$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge n\geq 2tk\}\,.$
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the ordered pairs in the modified recurrence:
$\gamma=\langle k,\,n\rangle,\,\beta_{0}=\langle k+1,\,n\rangle,\,\beta_{1}=\langle k,\,n-1\rangle,\,\dots,\,\beta_{t}=\langle k,\,n-t\rangle\,.$
The following three properties of the norm $\|\dots\|$ and the pairs $\gamma$, $\beta_{0}$, $\dots$, $\beta_{t}$ are easy to prove.
  1. 1.
    The norm $\|\dots\|$ attains a minimum value on the set $M$ (namely, zero).
  2. 2.
    If $\gamma\in M$ then $\beta_{i}\in M\cup N$ for every $i=0,1,\dots,t$.
  3. 3.
    For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every number $k\geq k_{0}$ has the required property. $\Box$
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j\in[t]$ ($\{1,2,\dots,t\}$), are $2t$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
$\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
We may assume that $\alpha_{j}\in\mathbb{N}$. We set $A:=\max(\{\alpha_{1},\dots,\alpha_{t}\})$ and define $u_{n}$ and $v_{n}$ as in Theorem 2.2. Then
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
and
$\sum_{n\geq 0}\frac{u_{n}}{n!}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
From the definition of $u_{n}$ and $A$ it easily follows that $u_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Since by the definition of $v_{n}$ we have
$|v_{n}|={\textstyle n!\cdot\big|\sum_{r=0}^{n}u_{r}/r!\big|=n!\cdot\big|\sum_{r=n+1}^{\infty}u_{r}/r!\big|}$
— now we invoke the contradictory assumption — we get the bound
$|v_{n}|$ $\leq$ ${\textstyle\frac{|u_{n+1}|}{n+1}+\frac{|u_{n+2}|}{(n+1)(n+2)}+\dots\leq cA^{n}\big(\frac{A}{n+1}+\frac{A^{2}}{(n+1)(n+2)}+\dots\big)}$
$\leq$ ${\textstyle cA^{n}\sum_{j\geq 0}\frac{A^{j}}{j!}=c\,\mathrm{e}^{A}\cdot A^{n}\,,}$
where $c\geq 0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{n\geq 0}(v_{n}-nv_{n-1})x^{n}=\sum_{n\geq 0}u_{n}x^{n}=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
(with $v_{-1}:=0$). However, the equality
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
We repeat that the proof in [2] is more general as it establishes the full LW theorem. In that proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of. The previous proof works with formal power series, but it is semiformal and not formal because the key player is the infinite sum
$\sum_{n\geq 0}\frac{u_{n}}{n!}=0$}
which is the value $f(1)$ of the formal power series $f(x)=\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}$. In fact, this is the only non-formal component in an otherwise formal proof.
It is clear that the previous proof satisfies condition (A), which states that all directly used sets are HMC. It works directly only with complex formal power series, which are maps from $\mathbb{N}_{0}$ to $\mathbb{C}$. As explained in Section 1, non-HMC sets like $\mathbb{C}[[x]]$ are not really needed in the proof. Condition (B), which states that no formalism of mathematical logic is used, is satisfied as well. Although the article [2] of Beukers, Bézivin, and Robba is a pioneering one, as it is the first AB proof of the transcendence of the number $\mathrm{e}$ (in fact, of course, of the full LW theorem), it is little known; in the database of Mathematical Reviews, it currently has only 11 citation (January 2026).
3 The proof by formal integration
We present the second AB &\& SF proof of the transcendence of $\mathrm{e}$. It is due to this author and first appeared in early versions of [7]. The idea is to translate Hilbert’s proof, which we described at the beginning, into the language of formal power series.
We work in the domains of formal power series $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$. We introduce further operations on $\mathbb{R}[[x]]$ and $\mathbb{R}\{x\}$ besides those we mentioned in Section 1. For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
$f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
The operation $f(x)\mapsto f(ax)$, which we call the constant inner product, is formal. Clearly, for every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
$(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
If $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ then $f(0x)=a_{0}$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
In the next proposition, whose straightforward proof we omit, we collect properties of the operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$.
Proposition 3.1
Let $f,g\in\mathbb{R}[[x]]$ and $a,b\in\mathbb{R}$. The following holds.
  1. 1.
    $(af+bg)^{\prime}=af^{\prime}+bg^{\prime}$, $(fg)^{\prime}=f^{\prime}g+fg^{\prime}$ and $f(ax)^{\prime}=af^{\prime}(ax)$.
  2. 2.
    $\int(af+bg)=a\int f+b\int g$ and $\int f(ax)=\frac{1}{a}(\int f)(ax)$ if $a\neq 0$.
  3. 3.
    If $f\in\mathbb{R}\{x\}$ then $f^{\prime},\int f\in\mathbb{R}\{x\}$.
  4. 4.
    $\big(\int f\big)^{\prime}=f$ and $\int(f^{\prime})=f-[x^{0}]f$.
We define the (formal) exponential.
Definition 3.2
The formal power series
$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
is called the (formal) exponential.
It is easy to show that $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and that $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a semiformal operation.
Definition 3.3 (semiformal shifts)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ be in $\mathbb{R}\{x\}$ and let $b\in\mathbb{R}$. We define the formal power series $f(x+b)$ in $\mathbb{R}[[x]]$ by
$f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and semiformal shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
In the next two propositions, we show that semiformal shifts behave in desirable ways.
Proposition 3.4
Let $f(x)\in\mathbb{R}\{x\}$ and $c,d\in\mathbb{R}$. Then
f((x+c)+d)=f(x+(c+d)).f((x+c)+d)=f(x+(c+d))\,.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $c,d$ be as stated, and let $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,f((x+c)+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus
${\textstyle[x^{n}]f((x+c)+d)=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
$={\textstyle\sum_{m\geq n}\sum_{l\geq m}\frac{l!}{n!(m-n)!(l-m)!}a_{l}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=}$
$=[x^{n}]f(x+(c+d))\,.$
In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
Proposition 3.5
Let $f(x),g(x)\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
$(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. The coefficient of $x^{n}$ in the formal power series $f(x+c)$, respectively $g(x+c)$, equals $\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}$, respectively $\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}$. Thus
${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
${\textstyle=\sum_{k=0}^{n}\sum_{m\geq k}\sum_{l\geq n-k}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=}$
$=[x^{n}](fg)(x+c)\,.$
In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([9]). $\Box$
The following semiformal exponential identity is somewhat counter-intuitive (at first we got it wrong).
Proposition 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$. Then
$\mathrm{Exp}\big((ax)+\frac{b}{a}\big)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$}
Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}=\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}$
$=$ $\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
$\Box$
We will use this identity only with a=1a=-1.
We introduce Newton integrals of convergent formal power series.
Definition 3.7 (Newton integral of fps)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
${\textstyle\int_{u}^{v}f=\int_{u}^{v}f(x)}$ $:=$ ${\textstyle\big(\int f\big)(v)-\big(\int f\big)(u)}$
$=$ ${\textstyle\sum_{n\geq 0}a_{n}\cdot\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}a_{n}\cdot\frac{u^{n+1}}{n+1}}$
of the values of the formal primitive of $f$ at the numbers $v$ and $u$.
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. The Newton integral is a semiformal operation because it uses two values of a formal power series. We prove three integral transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
Proposition 3.8
Let $f,g\in\mathbb{R}\{x\}$ and $u,v,a,b\in\mathbb{R}$. Then
$\int_{u}^{v}(af+bg)=a\int_{u}^{v}f+b\int_{u}^{v}g\,.$}
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in Proposition 3.1. $\Box$
Proposition 3.9
Let $f\in\mathbb{R}\{x\}$ and $u,v,w\in\mathbb{R}$. Then
$\int_{u}^{w}f=\int_{u}^{v}f+\int_{v}^{w}f\,.$}
Proof. Indeed,
$\big(\int f\big)(w)-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+\big(\int f\big)(w)-\big(\int f\big)(v)\,.$}
$\Box$
The third transformation needs some work.
Proposition 3.10
Let $f\in\mathbb{R}\{x\}$ and $u,v,b\in\mathbb{R}$. Then
$\int_{u}^{v}f(x+b)=\int_{u+b}^{v+b}f(x)\,.$}
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$, $v$ and $b$ be as stated. Then
${\textstyle\int_{u}^{v}f(x+b)=}$
${\textstyle=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{u^{n+1}}{n+1}}$
$={\textstyle\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}v^{n+1}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}u^{n+1}}$
${\textstyle=\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}v^{l}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}u^{l}}$
${\textstyle=\sum_{m\geq 0}a_{m}\frac{(v+b)^{m+1}}{m+1}-\sum_{m\geq 0}a_{m}\frac{(u+b)^{m+1}}{m+1}=\int_{u+b}^{v+b}f(x)\,.}$
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we add
$0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
In the fifth line, we use the binomial theorem. $\Box$
We introduce improper Newton integrals of convergent formal power series.
Definition 3.11 (improper Newton $\int$ of fps)
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the proper limit
$L=\lim_{n\to\infty}\int_{u}^{b_{n}}f\ \ (\in\mathbb{R})$}
exists, then $L$ does not depend on the sequence $(b_{n})$ and we define
$\int_{u}^{+\infty}f=\int_{u}^{+\infty}f(x):=L\,.$}
We call $L$ the improper Newton integral of $f(x)$ from $u$ to $\infty$.
We leave the proof of the independence of the limit $L$ on the sequences $(b_{n})$ to the reader as an exercise.
We prove the same three integral transformations as before.
Proposition 3.12
Let $f,g\in\mathbb{R}\{x\}$ and let $u,a,b\in\mathbb{R}$. Then the equality
$\int_{u}^{+\infty}(af+bg)=a\int_{u}^{+\infty}f+b\int_{u}^{+\infty}g$}
holds whenever the last two improper Newton integrals exist.
Proof. This follows from Proposition 3.8 and the arithmetic of limits of real sequences. $\Box$
Proposition 3.13
Let $f\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f=\int_{u}^{v}f+\int_{v}^{+\infty}f$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ exists. But then
${\textstyle\int_{u}^{+\infty}f=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
With the assumption of existence of $\int_{v}^{+\infty}f$ we argue similarly. $\Box$
The third transformation is immediate as well.
Proposition 3.14
Let $f(x)\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f(x+b)=\int_{u+b}^{+\infty}f(x)$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f(x)$ and $u,b$ be as stated. Let $(c_{n})$ be any real sequence going to $\infty$. We assume that the integral $I$ on the left-hand side exists, and prove that the integral on the right-hand side exists too and equals $I$. Indeed,
[…]
Similarly, if the integral on the right-hand side exists, then
[…]
$\Box$
We need to bound Newton integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)$ is in $\mathbb{Z}[x]$. We easily obtain the next bound based on two values of the primitive.
Proposition 3.15
For every numbers $i,k\in\mathbb{N}_{0}$,
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
Proof. The integrand
$\text{$\sum_{n\geq 0}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n\geq 0}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq\sum_{n\geq 0}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}\,.$}
$\Box$
It is crucial that the bound is exponential and not factorial in $k$.
The key step in Hilbert’s proof of the transcendence of $\mathrm{e}$ is Euler’s identity. We prove it in the language of improper Newton integrals of convergent formal power series.
Proposition 3.16
For every $k\in\mathbb{N}_{0}$, the integral
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n\geq k}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. Since
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,$,}
we have
$I_{0}=\lim_{n\to\infty}{\textstyle(1-\mathrm{Exp}(-a_{n}))-(1-\mathrm{Exp}(-0))=1-0=1\,.}$
Let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$, $k\in\mathbb{N}_{0}$, and let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}$
and $F_{k}=k\cdot F_{k-1}-F_{k}^{\prime}$. By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}-[x^{0}]F_{k}=F_{k}$. By Proposition 3.12 and induction,
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{k}}F_{k}^{\prime}}$
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}\big(F_{k}(a_{n})-F_{k}(0)\big)$
$=$ $k\cdot(k-1)!-0=k!\;.$
$\Box$
We are ready to translate Hilbert’s proof of the transcendence of $\mathrm{e}$ into formal power series. We again assume for the contradiction that
$a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral
$\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)\,$,}
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ is as in Section 1. This integral exists by Propositions 3.12 and 3.16. Using Proposition 3.13 we get that
$0$ $=$ $A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)+\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as at the beginning, only now they are defined in an AB &\& SF way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we calculate semiformally.
$B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
[…]
Using Propositions 3.12 and 3.16, and the definition of the polynomial $p_{r}(x)$, we obtain the properties of $B_{r}$ as before.
It remains to bound, in an AB &\& SF way, the numbers
$A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
The polynomial
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$
has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
$(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$
The previous proof is semiformal and not formal due to the use of various semiformal operations and many values of formal power series. It is clearly an AB proof.