### v6

# Two algebraic proofs of the transcendence of $\mathrm{e}$ based on formal power series

Abstract
We remind the classical analytical proof of the transcendence of $\mathrm{e}$ due to Hilbert. Using formal power series, we then give two algebraic proofs of this result. The first proof is a specialization of the proof of Beukers, Bézivin and Robba of the Lindemann–Weierstrass theorem. The second proof uses improper integrals of formal power series and is due to this author. We explain in what respect both proofs improve upon Hilbert’s proof.
1 Hilbert’s proof and how to improve it
The transcendence of the number $\mathrm{e}$ was first proven by Hermite [4]. Hilbert’s proof [5] goes as follows. We suppose for the contradiction that there exists $n+1$ integers $a_{0}$, $a_{1}$, $\dots$, $a_{n}$, not all of them 0, such that
$P:=a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0\,.$
We may assume that $a_{0}\neq 0$. For any $r\in\mathbb{N}$ ($\{1,2,\dots\}$) we define integral polynomials
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}\,$,
with $n$ as in $P$. We have the expansions
$p_{r}(x)=(-1)^{n(r+1)}(n!)^{r+1}\cdot x^{r}+ax^{r+1}+\dots$
and
$p_{r}(x+j)=bx^{r+1}+cx^{r+2}+\dots\,$,
where $j=1,2,\dots,n$ and $a,b,c,\dots$ are in $\mathbb{Z}$ (the integers). We consider improper integrals
$I_{r}=\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
Using the additivity of improper integrals, we express the product $PI_{r}$ as
$0$ $=$ $PI_{r}=A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x+\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}\,.$
As for the quantity $A_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{0}^{j}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x$, it is easy to estimate it by means of M-L bounds on integrals as
$|A_{r}|\leq c^{r},\ r\in\mathbb{N}\,$,
where $c\geq 1$ is a constant (depending only on the coefficients $a_{0}$, $a_{1}$, $\dots$, $a_{n}$).
We determine the second quantity
$B_{r}=\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x\,.$}
We claim that $B_{r}$ is an integral multiple of $r!$ and that $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. Since $c^{r}/r!\to 0$ as $r\to\infty$ for any constant $c\geq 1$, the above equalities
$A_{r}+B_{r}=0,\ r\in\mathbb{N}\,$,
are contradictory and we are done. We prove the claim by means of the integral identity that
$\int_{0}^{+\infty}x^{n}\mathrm{e}^{-x}\,\mathrm{d}x=n!$}
for every $n\in\mathbb{N}_{0}$ ($\{0,1,\dots\}$). The identity is due to Euler and is easy to establish by induction on $n$ via integration by parts. Using the linearity of improper integrals, we obtain for any integers $b_{0}$, $b_{1}$, $\dots$, $b_{m}$ with $m\in\mathbb{N}_{0}$ the more general identity that
$\int_{0}^{+\infty}\big(\sum_{j=0}^{m}b_{j}x^{j}\big)\mathrm{e}^{-x}\,\mathrm{d}x=\sum_{j=0}^{m}b_{j}\cdot j!\,.$}
Thus
$B_{r}$ $=$ ${\textstyle\sum_{j=0}^{n}a_{j}\mathrm{e}^{j}\int_{j}^{+\infty}p_{r}(x)\cdot\mathrm{e}^{-x}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{j}^{+\infty}p_{r}((x-j)+j)\cdot\mathrm{e}^{-(x-j)}\,\mathrm{d}x}$
$=$ ${\textstyle\sum_{j=0}^{n}a_{j}\int_{0}^{+\infty}p_{r}(y+j)\cdot\mathrm{e}^{-y}\,\mathrm{d}y}$
$=$ $\pm a_{0}(n!)^{r+1}\cdot r!+m_{r}\cdot(r+1)!\,$,
for some $m_{r}\in\mathbb{Z}$. In the second line, we used the identity that $\mathrm{e}^{a+b}=\mathrm{e}^{a}\mathrm{e}^{b}$. In the third line, we changed the variable to y=xjy=x-j. In the fourth line, we used the above generalization of Euler’s identity and the expansions of $p_{r}(x+j)$, $j=0,1,\dots,n$. It is clear that $B_{r}\in\mathbb{Z}$ and is divisible by $r!$. If $B_{r}=0$ then $(r+1)\,|\,a_{0}(n!)^{r+1}$. This divisibility is excluded when $r+1$ is coprime to the number $a_{0}\cdot n!$. Hence $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$. $\Box$
Baker [1, Theorem 1.2 on p. 4] gives a variant of the previous proof using only proper integrals $\int_{a}^{b}$. A similar proof appears in Conrad [3].
Hilbert’s proof is a mathematical gem, but there is still a large room for improvement. The functions $x^{n}\mathrm{e}^{-x}\colon[0,+\infty)\to\mathbb{R}$, which play a key role in the proof, are uncountable sets. Is it possible to give a proof of the transcendence of $\mathrm{e}$ without substantially using uncountable sets? (For example, when one proves the formula $\sum_{n\geq 0}a^{n}=\frac{1}{1-a}$ for $a\in(-1,1)$, one produces such a proof, even though the uncountable set (1,1)(-1,1) is mentioned; it is clear how to formulate (1,1)(-1,1) away. Of course, individual real numbers have to be represented by Dedekind’s cuts and not by Cantor’s equivalence classes of Cauchy sequences of fractions.) Yes, this is possible, such a proof was found already in 1990 by Beukers, Bézivin and Robba [2]; they actually proved in this way a much more general result, the Lindemann–Weierstrass theorem. What are the main countable players in their proof which replace the functions $x^{n}\mathrm{e}^{-x}$? The complex sequences
$(a_{n})=(a_{0},\,a_{1},\,a_{2},\,\dots),\ a_{n}\in\mathbb{C}\,$,
known also as formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with complex coefficients. In contrast to functions $x^{n}\mathrm{e}^{-x}$, the variable $x$ is now only formal.
Why do we hear almost nothing about [2]? (In March 2026 it had $1$ citation in Mathematical Reviews.) In the approach of eliminating uncountable sets from proofs, [2] is a remarkable achievement, even if the authors are probably not aware of it (that approach is not mentioned in [2]). However, only a small minority of professional mathematicians, to which this author belongs, care about the distinction of using or not using uncountable sets in proofs; see, for example, the forthcoming book [7].
The proof of Beukers, Bézivin and Robba deserves a new edition, so to say, and we review it in Section 2, but only in the specialization to the proof of the transcendence of $\mathrm{e}$. In Section 3 we present another uncountable-sets-free proof of the transcendence of $\mathrm{e}$ using formal power series. It was found by this author and is based on the natural idea of translating Hilbert’s proof to the language of formal power series.
To conclude the Introduction, we reiterate that the main contribution of the two proofs in the next two sections is not that they use formal power series; the main novelty is that they avoid (substantial use of) uncountable sets.
2 Specializing the proof of Beukers, Bézivin and Robba
The Lindemann–Weierstrass theorem says that
$b_{1}\mathrm{e}^{\alpha_{1}}+b_{2}\mathrm{e}^{\alpha_{2}}+\dots+b_{t}\mathrm{e}^{\alpha_{t}}\neq 0$
whenever $b_{j}$ and $\alpha_{j}$ are $2t\geq 2$ algebraic numbers in $\mathbb{C}$ such that every $b_{j}\neq 0$ and the $\alpha_{j}$ are mutually distinct. See Baker [1, Theorem 1.4 on pp. 6–8] for an analytic proof. Using formal power series, Beukers, Bézivin, and Robba [2] obtained another proof of the LW theorem. In this section, we specialize their proof to $b_{j},\alpha_{j}\in\mathbb{Z}$, which gives the transcendence of $\mathrm{e}$. In order to clearly present the logical structure of the proof, we reorganize the exposition in [2].
We first recall that a formal power series $f(x)$ in $\mathbb{C}[[x]]$ is rational if there exist two polynomials $p(x),q(x)\in\mathbb{C}[x]$ such that $q(0)=1$ and $q(x)f(x)=p(x)$; we may assume that $p(x)$ and $q(x)$ are coprime. If $f(x)$ and $g(x)$ in $\mathbb{C}[[x]]$ are rational, then so are $f(x)+g(x)$, $f(x)g(x)$, and $f^{\prime}(x)$. Any rational formal power series $f(x)\in\mathbb{C}[[x]]$ has a unique decomposition into partial fractions
$f(x)=r(x)+\sum_{j=1}^{t}\sum_{i=1}^{m_{j}}\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}\,$,}
where $r(x)\in\mathbb{C}[x]$, $t\in\mathbb{N}_{0}$, $m_{j}\in\mathbb{N}$, $\beta_{j,i},\alpha_{j}\in\mathbb{C}$ with $\beta_{j,m_{j}}\neq 0$, and the $\alpha_{j}$ are nonzero and mutually distinct. For $t=0$, the double sum is interpreted as 0, and the numbers $m_{j}$ and $\alpha_{j}$ come from the factorization
$q(x)=\prod_{j=1}^{t}(1-\alpha_{j}x)^{m_{j}}\,.$}
The partial fractions are the formal power series
$\frac{\beta_{j,i}}{(1-\alpha_{j}x)^{i}}=\sum_{n\geq 0}\binom{-i}{n}\beta_{j,i}\alpha_{j}^{n}\cdot x^{n}\,.$}
We say that $1/\alpha_{j}$ is a pole of $f(x)$ with order $m_{j}$.
Proposition 2.1
If $p(x)\in\mathbb{C}[x]$, $m\in\mathbb{N}_{0}$, $c\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$) and if $A(x)$ in $\mathbb{C}[[x]]$ is rational, then the rational formal power series
$p(x)A(x)+cx^{m}A^{\prime}(x)$
either has no pole (is a polynomial) or all its poles have orders at least $2$.
Proof. The product $p(x)A(x)$ only possibly decreases the orders of the poles of $A(x)$ or cancels them and does not create any new poles. The product and derivative $cx^{m}A^{\prime}(x)$ preserve all poles of $A(x)$, increase their orders by $1$, and do not create any new poles. The sum $p(x)A(x)+cx^{m}A^{\prime}(x)$ merges the two sets of poles. It follows that the resulting set of poles with their orders is that of $cx^{m}A^{\prime}(x)$. $\Box$
We extracted the following theorem on rational formal power series from the arguments in [2].
Theorem 2.2
Let $t\in\mathbb{N}$, $b_{1}$, $b_{2}$, $\dots$, $b_{t}$ be nonzero integers and $\alpha_{1}$, $\alpha_{2}$, $\dots$, $\alpha_{t}$ be distinct natural numbers. For every $n\in\mathbb{N}_{0}$ we define integers
$u_{n}=b_{1}\alpha_{1}^{n}+\dots+b_{t}\alpha_{t}^{n}\,\text{ and }\,v_{n}=n!\cdot{\textstyle\sum_{r=0}^{n}u_{r}/r!}\,.$
Suppose that for some constant $A\geq 1$ we have bound
$v_{n}=O(A^{n})\,\text{ for }\,n\in\mathbb{N}_{0}\,.$
Then the formal power series $\sum_{n\geq 0}v_{n}x^{n}$ is rational, because
$\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\in\mathbb{Z}[x]$}
for some $k\in\mathbb{N}$, where the integers $a_{j}$ are the coefficients in
$\prod_{j=1}^{t}(1-\alpha_{j}x)=1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\,.$}
Proof. For $k,n\in\mathbb{N}_{0}$ we define integers $v_{n}(k)$ by the relations
$\sum_{n\geq 0}v_{n}(k)x^{n}=\big(1-a_{1}x-a_{2}x^{2}-\dots-a_{t}x^{t}\big)^{k}\sum_{n\geq 0}v_{n}x^{n}\,.$}
Then for every $n\geq t$ and $k\in\mathbb{N}_{0}$ we have the recurrence
$v_{n}(k+1)=v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)\,.$
We establish for every $n,k\in\mathbb{N}_{0}$ with $n\geq tk$ two properties of the numbers $v_{n}(k)$.
  1. 1.
    We have the bound $|v_{n}(k)|\leq cA^{n}C^{k}$ where $C=1+|a_{1}|+\dots+|a_{t}|$ and $c\geq 0$ is a constant.
  2. 2.
    $k!\,|\,v_{n}(k)$.
The bound in item 1 follows by induction on $k$ from the assumed bound on $v_{n}=v_{n}(0)$. For $k=0$ the bound holds because it is the assumed bound on $v_{n}$. Let $n,k\in\mathbb{N}_{0}$ with $n\geq t(k+1)$. Then $n-t\geq tk$, and by the above recurrence and induction we have
$|v_{n}(k+1)|$ $=$ $|v_{n}(k)-a_{1}v_{n-1}(k)-\dots-a_{t}v_{n-t}(k)|$
$\leq$ $c\big(A^{n}C^{k}+A^{n-1}C^{k}|a_{1}|+\dots+A^{n-t}C^{k}|a_{t}|\big)\leq cA^{n}C^{k+1}\,.$
In order to prove the divisibility in item 2, we note that $v_{0}=u_{0}$ and that for $n\in\mathbb{N}$,
$\frac{v_{n}}{n!}-\frac{v_{n-1}}{(n-1)!}=\frac{u_{n}}{n!}\,\text{ (hence $v_{n}-nv_{n-1}=u_{n}$)}\,.$}
For $k=0$ the divisibility trivially holds. So let $k\in\mathbb{N}$. We set $(0)_{0}=(n)_{0}=1$ and $(n)_{m}=n(n-1)\dots(n-m+1)$ for $m\in\mathbb{N}$. For $n\in\mathbb{N}_{0}$ we combine $\min(n,k)$ displayed differences, multiply the result by $n!$, and get that
$v_{n}=(n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}+(n)_{k}v_{n-k}$}\,,
where the $u_{\dots}$ and $v_{\dots}$ with negative indices are defined arbitrarily, for example as 0. We denote the last summand $(n)_{k}v_{n-k}$ ($\in\mathbb{Z}$) by $w_{n}$. Note that always $k!\,|\,w_{n}$ because $w_{n}=k!\cdot\binom{n}{k}v_{n-k}$ for $n\geq k$, and $w_{n}=0$ for $0\leq n<k$. The difference of the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ and $W(x)=\sum_{n\geq 0}w_{n}x^{n}$ is
$V(x)-W(x)=\sum_{n\geq 0}\big((n)_{0}u_{n}+(n)_{1}u_{n-1}+\dots+(n)_{k-1}u_{n-k+1}\big)x^{n}\,.$}
We claim that
$V(x)-W(x)=\frac{p(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{k}}=:\frac{p(x)}{q(x)}$}
for some polynomial $p(x)\in\mathbb{Z}[x]$ with degree less than $tk$. Then for every $n\geq tk$ we have
$v_{n}(k)=[x^{n}]q(x)V(x)=[x^{n}]q(x)W(x)$
and $k!\,|\,v_{n}(k)$ because every coefficient in $W(x)$ is divisible by $k!$. We prove the claim: for every $r$ with $0\leq r\leq k-1$ we indeed have, since $\binom{n}{r}=(-1)^{n-r}\binom{-r-1}{n-r}$ for $n\geq r$, that
${\textstyle\sum_{n\geq 0}(n)_{r}u_{n-r}x^{n}}$ $=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}\sum_{n\geq r}\binom{n}{r}\alpha_{j}^{n-r}x^{n-r}}$
$=$ ${\textstyle\sum_{j=1}^{t}r!\cdot b_{j}x^{r}(1-\alpha_{j}x)^{-r-1}=\frac{p_{r}(x)}{(1-a_{1}x-\dots-a_{t}x^{t})^{r+1}}}$
for some polynomial $p_{r}(x)\in\mathbb{Z}[x]$ with degree less than $t(r+1)$.
Now the two properties of numbers $v_{n}(k)$ imply that if $n\geq tk$ and $v_{n}(k)\neq 0$, then
$k!\leq|v_{n}(k)|\leq cA^{n}C^{k}\,.$
In other words, if $n\geq tk$ and if $k!>cA^{n}C^{k}$, then $v_{n}(k)=0$. Using this implication and the modification
$v_{n}(k)=v_{n}(k+1)+a_{1}v_{n-1}(k)+\dots+a_{t}v_{n-t}(k),\ n\geq t\,$,
of the above recurrence, we show that there exists a $k\in\mathbb{N}$ such that $v_{n}(k)=0$ for every large $n$. This will prove that the formal power series $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ is rational. So let $k_{0}\in\mathbb{N}$ be such that $k!>cA^{2tk}C^{k}$ for every $k\geq k_{0}$, and let
$N$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge tk\leq n\leq 2tk\}\,\text{ and}$
$M$ $=$ $\{\langle k,\,n\rangle\in\mathbb{N}^{2}\colon\;k\geq k_{0}\wedge n\geq 2tk\}\,.$
Thus $\langle k,n\rangle\in N$ $\Rightarrow$ $v_{n}(k)=0$. For $\langle k,n\rangle\in\mathbb{N}^{2}$ we consider the norm $\|\langle k,n\rangle\|=n-2tk$ ($\in\mathbb{Z}$). We name the ordered pairs in the modified recurrence:
$\gamma=\langle k,\,n\rangle,\,\beta_{0}=\langle k+1,\,n\rangle,\,\beta_{1}=\langle k,\,n-1\rangle,\,\dots,\,\beta_{t}=\langle k,\,n-t\rangle\,.$
The following three properties of the norm $\|\dots\|$ and the pairs $\gamma$, $\beta_{0}$, $\dots$, $\beta_{t}$ are easy to prove.
  1. 1.
    The norm $\|\dots\|$ attains a minimum value on the set $M$ (namely, zero).
  2. 2.
    If $\gamma\in M$ then $\beta_{i}\in M\cup N$ for every $i=0,1,\dots,t$.
  3. 3.
    For every $i=0,1,\dots,t$, the norm $\|\beta_{i}\|<\|\gamma\|$.
From these three properties, from the property of the set $N$ and from the modified recurrence, it follows by induction on the norm that $v_{n}(k)=0$ for every $\langle k,n\rangle\in M$. This concludes the proof because we see that every number $k\geq k_{0}$ has the required property. $\Box$
We prove, with the help of Proposition 2.1 and Theorem 2.2, that the number $\mathrm{e}$ is transcendental. We assume, for the contradiction, that $b_{j},\alpha_{j}\in\mathbb{Z}$, $j=1,2,\dots,t$, are $2t\geq 2$ integers such that every $b_{j}\neq 0$, the $\alpha_{j}$ are mutually distinct, and that
$\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
We may assume that $\alpha_{j}\in\mathbb{N}$. We set $A:=\max(\{\alpha_{1},\dots,\alpha_{t}\})$ and define $u_{n}$ and $v_{n}$ as in Theorem 2.2. Then
$\sum_{n\geq 0}\frac{u_{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\sum_{n\geq 0}\frac{\alpha_{j}^{n}}{n!}x^{n}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}x}$}
and
$\sum_{n\geq 0}\frac{u_{n}}{n!}=\sum_{j=1}^{t}b_{j}\mathrm{e}^{\alpha_{j}}=0\,.$}
From the definition of $u_{n}$ and $A$ it easily follows that $u_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Since by the definition of $v_{n}$ we have
$|v_{n}|={\textstyle n!\cdot\big|\sum_{r=0}^{n}u_{r}/r!\big|=n!\cdot\big|\sum_{r=n+1}^{\infty}u_{r}/r!\big|}$
— now we invoke the contradictory assumption — we get the bound
$|v_{n}|$ $\leq$ ${\textstyle\frac{|u_{n+1}|}{n+1}+\frac{|u_{n+2}|}{(n+1)(n+2)}+\dots\leq cA^{n}\big(\frac{A}{n+1}+\frac{A^{2}}{(n+1)(n+2)}+\dots\big)}$
$\leq$ ${\textstyle cA^{n}\sum_{j\geq 0}\frac{A^{j}}{j!}=c\,\mathrm{e}^{A}\cdot A^{n}\,,}$
where $c\geq 0$ is an absolute constant. Thus $v_{n}=O(A^{n})$ for $n\in\mathbb{N}_{0}$. Let $V(x)=\sum_{n\geq 0}v_{n}x^{n}$ ($\in\mathbb{Z}[[x]]$). As we know from the previous proof, $v_{n}-nv_{n-1}=u_{n}$ for $n\in\mathbb{N}$ and $v_{0}=u_{0}$. In terms of formal power series we get
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{n\geq 0}(v_{n}-nv_{n-1})x^{n}=\sum_{n\geq 0}u_{n}x^{n}=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
(with $v_{-1}:=0$). However, the equality
$(1-x)V(x)-x^{2}V^{\prime}(x)=\sum_{j=1}^{t}\frac{b_{j}}{1-\alpha_{j}x}$}
contradicts Proposition 2.1 because by Theorem 2.2 the formal power series $V(x)$ is rational. $\Box$
In the original proof, one can only achieve that $u_{n},v_{n}\in\mathbb{Q}$, and denominators have to be taken care of.
3 The proof by semiformal integration
We work in the domains of formal power series $\mathbb{R}[[x]]$ and convergent formal power series $\mathbb{R}\{x\}$; the latter are formal power series with an infinite radius of convergence. For $f(x)\in\mathbb{R}[[x]]$ and $n\in\mathbb{N}_{0}$ we denote by $x^{n}]f(x)$ the coefficient of $x^{n}$ in $f(x)$.
For any $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ in $\mathbb{R}[[x]]$ and any $a\in\mathbb{R}$ we define the formal power series
$f(ax)=\sum_{n\geq 0}a^{n}a_{n}x^{n}\,.$}
The family of operations
$\{f(x)\mapsto f(ax)\colon\;a\in\mathbb{R}\}\,$,
which we call constant inner products, consists of formal unary operations. For every $a,b,c\in\mathbb{R}$ and every $f(x)$ and $g(x)$ in $\mathbb{R}[[x]]$ we have $f(1x)=f(x)$,
$(af+bg)(cx)=af(cx)+bg(cx)\,\text{ and }\,(fg)(cx)=f(cx)g(cx)\,.$
Also, $f(0x)=[x^{0}]f$. If $f(x)\in\mathbb{R}\{x\}$ then also $f(cx)\in\mathbb{R}\{x\}$. By f(x)f(-x) we mean f((1)x)f((-1)x).
In the next proposition, whose straightforward proof we omit, we collect properties of the unary operations $f\mapsto f^{\prime}$ and $f\mapsto\int f$ on $\mathbb{R}[[x]]$. Recall that
$\int\sum_{n\geq 0}a_{n}x^{n}=\sum_{n\geq 0}\frac{a_{n}}{n+1}x^{n+1}=\sum_{n\geq 1}\frac{a_{n-1}}{n}x^{n}\,.$}
Proposition 3.1
Let $f,g\in\mathbb{R}[[x]]$ and $a,b\in\mathbb{R}$. The following holds.
  1. 1.
    $(af+bg)^{\prime}=af^{\prime}+bg^{\prime}$, $(fg)^{\prime}=f^{\prime}g+fg^{\prime}$ and $f(ax)^{\prime}=af^{\prime}(ax)$.
  2. 2.
    $\int(af+bg)=a\int f+b\int g$ and $\int f(ax)=\frac{1}{a}(\int f)(ax)$ if $a\neq 0$.
  3. 3.
    If $f\in\mathbb{R}\{x\}$ then $f^{\prime},\int f\in\mathbb{R}\{x\}$.
  4. 4.
    $\big(\int f\big)^{\prime}=f$ and $\int(f^{\prime})=f-[x^{0}]f$.
We define the (formal) exponential.
Definition 3.2
The formal power series
$\mathrm{Exp}(x)=\sum_{n\geq 0}\frac{1}{n!}x^{n}=1+x+\frac{1}{2}x^{2}+\frac{1}{6}x^{3}+\dots\ \ (\in\mathbb{Q}[[x]])$}
is called the (formal) exponential.
Clearly, $\mathrm{Exp}\in\mathbb{R}\{x\}$, $\mathrm{Exp}^{\prime}=\mathrm{Exp}$ and $\int\mathrm{Exp}=\mathrm{Exp}-1$.
We introduce a family of semiformal unary operations.
Definition 3.3 (shifts)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ be in $\mathbb{R}\{x\}$ and let $b\in\mathbb{R}$. We define the formal power series $f(x+b)$ in $\mathbb{R}[[x]]$ by
$f(x+b)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)x^{n}\;.$}
We call these operations on $\mathbb{R}[[x]]$ semiformal because coefficients in them arise by limit transitions; in the standard formal operations they are obtained by finite expressions. The coefficients in $f(x+b)$ are correctly defined because $f(x)$ is convergent. It is easy to show that $f(x+b)\in\mathbb{R}\{x\}$. Clearly, $f(x+0)=f(x)$ and if $f(x)\in\mathbb{R}[x]$ then $f(x)\mapsto f(x+b)$ is the usual polynomial substitution. We combine the operations of constant inner product and shift by means of brackets. For example, $f((x+a)+b)$ is $g(x+b)$ where $g(x)=f(x+a)$ or $f((ax)+b)$ is $g(x+b)$ where $g(x)=f(ax)$ or $f(a(x+b))$ is $g(ax)$ where $g(x)=f(x+b)$ and so on.
We show that shifts behave in desirable ways.
Proposition 3.4
Let $f(x)\in\mathbb{R}\{x\}$ and $c,d\in\mathbb{R}$. Then
f((x+c)+d)=f(x+(c+d)).f((x+c)+d)=f(x+(c+d))\,.
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$ and $c,d$ be as stated, and let $f(x+c)=\sum_{n\geq 0}b_{n}x^{n}$. Then
$b_{n}=\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,f((x+c)+d)=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}b_{m}d^{m-n}\big)x^{n}\,.$}
Thus
${\textstyle[x^{n}]f((x+c)+d)=\sum_{m\geq n}\binom{m}{n}\big(\sum_{l\geq m}\binom{l}{m}a_{l}c^{l-m}\big)d^{m-n}=}$
$={\textstyle\sum_{m\geq n}\sum_{l\geq m}\frac{l!}{n!(m-n)!(l-m)!}a_{l}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{m=n}^{l}\binom{l-n}{l-m}c^{l-m}d^{m-n}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}\sum_{j=0}^{l-n}\binom{l-n}{j}c^{j}d^{l-n-j}=}$
${\textstyle=\sum_{l\geq n}\binom{l}{n}a_{l}(c+d)^{l-n}=[x^{n}]f(x+(c+d))\,.}$
In the third line, we changed the order of summation. This is allowed because $f(x)$ is convergent. In the fourth line, we used the binomial theorem. $\Box$
Proposition 3.5
Let $f(x),g(x)\in\mathbb{R}\{x\}$ and $c\in\mathbb{R}$. Then
$(f\cdot g)(x+c)=f(x+c)\cdot g(x+c)\;.$
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $g(x)=\sum_{n\geq 0}b_{n}x^{n}$ and $c$ be as stated. Then the coefficients $x^{n}]f(x+c)$ and $x^{n}]g(x+c)$ equal to
$\sum_{m\geq n}\binom{m}{n}a_{m}c^{m-n}\,\text{ and }\,\sum_{l\geq n}\binom{l}{n}b_{l}c^{l-n}\,$,}
respectively. Thus
${\textstyle[x^{n}]f(x+c)g(x+c)=\sum_{k=0}^{n}\big(\sum_{m\geq k}\binom{m}{k}a_{m}c^{m-k}\sum_{l\geq n-k}\binom{l}{n-k}b_{l}c^{l+k-n}\big)}$
${\textstyle=\sum_{k=0}^{n}\sum_{m\geq k}\sum_{l\geq n-k}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\sum_{k=0}^{n}\binom{m}{k}\binom{l}{n-k}a_{m}b_{l}c^{l+m-n}=}$
${\textstyle=\sum_{\begin{subarray}{c}m,\,l\geq 0\\ m+l\geq n\end{subarray}}\binom{m+l}{n}a_{m}b_{l}c^{l+m-n}}$
${\textstyle=\sum_{j\geq n}\binom{j}{n}\big(\sum_{m=0}^{j}a_{m}b_{j-m}\big)c^{j-n}=[x^{n}](fg)(x+c)\,.}$
In the third line, we changed the order of summation. This is allowed since $f(x)$ is convergent. In the fourth line, we used Vandermonde’s identity ([6]). $\Box$
The following semiformal exponential identity is perhaps counter-intuitive; at first we got it wrong.
Proposition 3.6
Let $a,b\in\mathbb{R}$ with $a\neq 0$. Then
$\mathrm{Exp}\big((ax)+\frac{b}{a}\big)=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$}
Proof. Let $a$, $b$ be as stated, and let $f(x)=\mathrm{Exp}(ax)$. Then
${\textstyle f\big(x+\frac{b}{a}\big)}$ $=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}(a^{m}/m!)(b/a)^{m-n}\big)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\sum_{m\geq n}\frac{b^{m-n}}{(m-n)!}(a^{n}/n!)x^{n}}$
$=$ ${\textstyle\sum_{n\geq 0}\big(\sum_{l\geq 0}\frac{b^{l}}{l!}\big)(a^{n}/n!)x^{n}}=\mathrm{e}^{b}\cdot\mathrm{Exp}(ax)\,.$
$\Box$
We use this identity only for a=1a=-1.
We introduce a family of semiformal functionals on $\mathbb{R}[[x]]$.
Definition 3.7 (Newton integral of fps)
Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}\in\mathbb{R}\{x\}$ and let $u,v\in\mathbb{R}$. We define the Newton integral of $f(x)$ from $u$ to $v$ as the difference
${\textstyle\int_{u}^{v}f=\int_{u}^{v}f(x)}$ $:=$ ${\textstyle\big(\int f\big)(v)-\big(\int f\big)(u)}$
$=$ ${\textstyle\sum_{n\geq 0}a_{n}\cdot\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}a_{n}\cdot\frac{u^{n+1}}{n+1}}$
of the values of the formal primitive of $f$ at the numbers $v$ and $u$.
By item 3 of Proposition 3.1, $\int f\in\mathbb{R}\{x\}$ and the integral is always defined. We obtain three transformations for it: linear combination, additive decomposition, and shift of the integration interval. The first two transformations are straightforward.
Proposition 3.8
Let $f,g\in\mathbb{R}\{x\}$ and $u,v,a,b\in\mathbb{R}$. Then
$\int_{u}^{v}(af+bg)=a\int_{u}^{v}f+b\int_{u}^{v}g\,.$}
Proof. This follows from the equality $\int(af+bg)=a\int f+b\int g$ in item 2 of Proposition 3.1. $\Box$
Proposition 3.9
Let $f\in\mathbb{R}\{x\}$ and $u,v,w\in\mathbb{R}$. Then
$\int_{u}^{w}f=\int_{u}^{v}f+\int_{v}^{w}f\,.$}
Proof. Indeed,
$\big(\int f\big)(w)-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+\big(\int f\big)(w)-\big(\int f\big)(v)\,.$}
$\Box$
The third transformation needs some work.
Proposition 3.10
Let $f\in\mathbb{R}\{x\}$ and $u,v,b\in\mathbb{R}$. Then
$\int_{u}^{v}f(x+b)=\int_{u+b}^{v+b}f(x)\,.$}
Proof. Let $f(x)=\sum_{n\geq 0}a_{n}x^{n}$, $u$, $v$ and $b$ be as stated. Then
${\textstyle\int_{u}^{v}f(x+b)=}$
${\textstyle=\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{v^{n+1}}{n+1}-\sum_{n\geq 0}\big(\sum_{m\geq n}\binom{m}{n}a_{m}b^{m-n}\big)\frac{u^{n+1}}{n+1}}$
$={\textstyle\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}v^{n+1}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{n=0}^{m}\binom{m+1}{n+1}b^{m-n}u^{n+1}}$
${\textstyle=\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}v^{l}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\sum_{l=0}^{m+1}\binom{m+1}{l}b^{m+1-l}u^{l}}$
${\textstyle=\sum_{m\geq 0}a_{m}\frac{(v+b)^{m+1}}{m+1}-\sum_{m\geq 0}a_{m}\frac{(u+b)^{m+1}}{m+1}=\int_{u+b}^{v+b}f(x)\,.}$
In the third line, we used the identity $\binom{m+1}{n+1}=\frac{m+1}{n+1}\binom{m}{n}$ and we exchanged the order of summation. This is allowed since the involved series absolutely converge. In the fourth line, we added
$0=\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}v^{0}-\sum_{m\geq 0}\frac{a_{m}}{m+1}\binom{m+1}{0}b^{m+1}u^{0}\,.$}
In the fifth line, we used the binomial theorem. $\Box$
We introduce another family of semiformal functionals on $\mathbb{R}[[x]]$.
Definition 3.11 (improper Newton $\int$ of fps)
Let $u\in\mathbb{R}$ and $f(x)\in\mathbb{R}\{x\}$. If for every real sequence $(b_{n})$ going to $\infty$ the proper limit
$L=\lim_{n\to\infty}\int_{u}^{b_{n}}f\ \ (\in\mathbb{R})$}
exists, then $L$ does not depend on the sequence $(b_{n})$ and we define
$\int_{u}^{+\infty}f=\int_{u}^{+\infty}f(x):=L\,.$}
We call $L$ the improper Newton integral of $f(x)$ from $u$ to $\infty$.
We leave the proof of the independence of the limit $L$ on the sequence $(b_{n})$ to the reader as an exercise.
We obtain for the improper integral the same three integral transformations as before.
Proposition 3.12
Let $f,g\in\mathbb{R}\{x\}$ and let $u,a,b\in\mathbb{R}$. Then the equality
$\int_{u}^{+\infty}(af+bg)=a\int_{u}^{+\infty}f+b\int_{u}^{+\infty}g$}
holds whenever the last two improper Newton integrals exist.
Proof. This follows from Proposition 3.8 and the arithmetic of limits of real sequences. $\Box$
Proposition 3.13
Let $f\in\mathbb{R}\{x\}$ and $u,v\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f=\int_{u}^{v}f+\int_{v}^{+\infty}f$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f$, $u$ and $v$ be as stated, let $(b_{n})$ be any real sequence going to $\infty$ and let us suppose that the former improper integral exists. Thus the limit $L=\lim_{n\to\infty}\big(\int f\big)(b_{n})$ ($\in\mathbb{R}$) exists. But then
${\textstyle\int_{u}^{+\infty}f=L-\big(\int f\big)(u)=\big(\int f\big)(v)-\big(\int f\big)(u)+L-\big(\int f\big)(v)=}$
${\textstyle=\int_{u}^{v}f+\int_{v}^{+\infty}f\,.}$
In the other way we argue similarly. $\Box$
The third transformation is immediate as well.
Proposition 3.14
Let $f(x)\in\mathbb{R}\{x\}$ and $u,b\in\mathbb{R}$. The equality
$\int_{u}^{+\infty}f(x+b)=\int_{u+b}^{+\infty}f(x)$}
holds whenever one of the two improper Newton integrals exists.
Proof. Let $f(x)$ and $u,b$ be as stated. Let $(c_{n})$ be any real sequence going to $\infty$. We assume that the integral $I$ on the left-hand side exists, and prove that the integral on the right-hand side exists too and equals $I$. Indeed,
[…]
Similarly, if the integral on the right-hand side exists, then
[…]
$\Box$
We need bounds on integrals $\int_{0}^{i}p(x)\mathrm{Exp}(-x)$ where $i\in\mathbb{N}_{0}$ and $p(x)\in\mathbb{Z}[x]$. We obtain them via the next bound which is based on two values of the primitive.
Proposition 3.15
For every numbers $i,k\in\mathbb{N}_{0}$,
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq i^{k+1}\mathrm{e}^{i}\,.$}
Proof. The integrand
$\text{$\sum_{n\geq 0}\frac{(-1)^{n}}{n!}x^{n+k}$ has the primitive $\sum_{n\geq 0}\frac{(-1)^{n}}{n!(n+k+1)}x^{n+k+1}$}\,.$}
Thus
$\big|\int_{0}^{i}x^{k}\cdot\mathrm{Exp}(-x)\big|\leq\sum_{n\geq 0}\frac{1}{n!}i^{n+k+1}+0=i^{k+1}\mathrm{e}^{i}\,.$}
$\Box$
As we know, it is crucial that the bound is exponential and not factorial in $k$.
We prove Euler’s identity without using uncountable sets.
Proposition 3.16 (Euler’s identity)
For every $k\in\mathbb{N}_{0}$, the integral
$I_{k}=\int_{0}^{+\infty}x^{k}\cdot\mathrm{Exp}(-x)=\int_{0}^{+\infty}\sum_{n\geq k}\frac{(-1)^{n-k}}{(n-k)!}x^{n}=k!\,.$}
Proof. We proceed by induction on $k$. Let $(a_{n})$ be any real sequence going to $\infty$. Since
$\int\mathrm{Exp}(-x)=\sum_{n\geq 0}\frac{(-1)^{n}}{(n+1)!}x^{n+1}=1-\mathrm{Exp}(-x)\,$,}
we have
$I_{0}=\lim_{n\to\infty}{\textstyle(1-\mathrm{Exp}(-a_{n}))-(1-\mathrm{Exp}(-0))=1-0=1\,.}$
Let $F_{k}=x^{k}\cdot\mathrm{Exp}(-x)$, $k\in\mathbb{N}_{0}$, and let $k\in\mathbb{N}$. By item 1 of Proposition 3.1 and the derivative of $\mathrm{Exp}(-x)$,
$F_{k}^{\prime}=\big(x^{k}\cdot\mathrm{Exp}(-x)\big)^{\prime}=kx^{k-1}\cdot\mathrm{Exp}(-x)-x^{k}\cdot\mathrm{Exp}(-x)=k\cdot F_{k-1}-F_{k}$
and $F_{k}=k\cdot F_{k-1}-F_{k}^{\prime}$. By item 4 of Proposition 3.1, $\int F_{k}^{\prime}=F_{k}-[x^{0}]F_{k}=F_{k}$. By Proposition 3.12 and induction,
$I_{k}$ $=$ $\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}}=k\cdot\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k-1}-}\lim_{n\to\infty}{\textstyle\int_{0}^{a_{n}}F_{k}^{\prime}}$
$=$ $k\cdot I_{k-1}-\lim_{n\to\infty}\big(F_{k}(a_{n})-F_{k}(0)\big)$
$=$ $k\cdot(k-1)!-0=k!\;.$
$\Box$
We are ready to state Hilbert’s proof in an uncountable-sets-free way. We again assume for the contradiction that
$a_{0}+a_{1}\mathrm{e}+a_{2}\mathrm{e}^{2}+\dots+a_{n}\mathrm{e}^{n}=0$
where $a_{i}\in\mathbb{Z}$, $n\in\mathbb{N}$, and $a_{0}\neq 0$. We multiply this equality by the improper Newton integral
$\int_{0}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)\,$,}
where $p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$ is as in Section 1. This integral exists by Propositions 3.12 and 3.16. Using Proposition 3.13 we get that
$0$ $=$ $A_{r}+B_{r}$
$:=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)+\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\cdot\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
for every $r\in\mathbb{N}$. The numbers $A_{r},B_{r}\in\mathbb{R}$ are the same as before, only now they are defined in an uncountable-sets-free way. We get the same contradiction as in Section 1: every number $B_{r}$ is an integral multiple of $r!$, $B_{r}\neq 0$ for infinitely many $r\in\mathbb{N}$, and numbers $A_{r}$ are exponentially bounded. The computation of $B_{r}$ proceeds along the same lines as in Section 1, the only difference is that now we compute semiformally.
$B_{r}$ $=$ ${\textstyle\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{i}^{+\infty}p_{r}(x)\cdot\mathrm{Exp}(-x)}$
[…]
Using Propositions 3.12 and 3.16, and the definition of $p_{r}(x)$, we obtain the above properties of $B_{r}$.
It remains to bound, in an uncountable-sets-free way, the numbers
$A_{r}=\sum_{i=0}^{n}a_{i}\mathrm{e}^{i}\int_{0}^{i}p_{r}(x)\cdot\mathrm{Exp}(-x)\,.$}
The polynomial
$p_{r}(x)=x^{r}((x-1)(x-2)\dots(x-n))^{r+1}$
has degree less than $(n+1)(r+1)$ and each coefficient in it is in absolute value at most $l^{r+1}$ where $l\in\mathbb{N}$ depends only on $n$. By Propositions 3.8 and 3.15 each of the above $n+1$ Newton integrals in $A_{r}$ is in absolute value at most
$(n+1)(r+1)\cdot l^{r+1}\cdot n^{(n+1)(r+1)}\cdot\mathrm{e}^{n}\,.$
Thus $|A_{r}|\leq c^{r}$ for every $r\in\mathbb{N}$ and some constant $c\geq 1$. We get a contradiction as before. $\Box$